### v1

# Incorporate LLMs with Influential Recommender System

Abstract.
Recommender systems have achieved increasing accuracy over the years. However, this precision often leads users to narrow their interests, resulting in issues such as limited diversity and the creation of echo chambers. Current research addresses these challenges through proactive recommender systems by recommending a sequence of items (called influence path) to guide user interest to the target item. However, existing methods struggle to construct a coherent influence path that build up with items the user is likely to enjoy. In this paper, we leverage the Large Language Model’s (LLMs) exceptional ability for path planning and instruction following, introducing a novel approach named LLM-based Influence Path Planing (LLM-IPP). Our approach maintains coherence between consecutive recommendations, and enhances user acceptability to the recommended items. To evaluate LLM-IPP, we implement various user simulators and metrics to measure user acceptability and path coherence. Experimental results demonstrate that LLM-IPP significantly outperforms traditional proactive recommender systems. This study is pioneering in integrating LLMs into proactive recommender systems, offering a reliable and user-engaging methodology for future recommendation technologies.
Proactive Recommender Systems, Large Language Model, Influence Path, User Simulator
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXbooktitle: ccs: Information systems Recommender systems
1. Introduction
Figure 1. Illustration of Passive and Proactive Recommendation.
Recommender systems play an important role in enhancing user experiences. Traditionally, recommender systems infer user interest from historical behavior patterns and recommend items aligned with the user’s interest. However, it will narrow the user’s interest and cause the echo chamber phenomenon by continually catering to historical interest (Gao et al., 2023b, a). To avoid this problem, researchers have proposed the task of proactive recommender system (Zhu et al., 2023), which aims to actively guide the user interest with a sequence of items (called influence path) toward a given target item, as shown in Figure 1. The influence path gradually guides shifts in user interests through continuous recommendations, facilitating the discovery of new interests. Therefore, influence path planning is crucial for enhancing the efficacy of proactive recommender systems in guiding user interests.
The most pertinent work is Influential Recommender System (IRS) (Zhu et al., 2023), which frames proactive recommendation as a sequential recommendation problem and constructs the influence path using a Transformer-based sequential model. Such a sequential model learns the interest-shifting patterns from users’ historical data and encodes the target item’s characteristics to generate the influence path. However, IRS encounters several critical issues: 1) There is a low probability of the target item being included in the influence path, with a success rate below 30% on MovieLens-1M (refer to Table 1); and 2) the adjacent items in the path are not sufficiently coherent, hindering the effective guidance of user interest. Moreover, directly incorporating the target item into the influence path to address the first issue will further diminish the path’s coherence.
Inspired by these, the key objectives of influence path planning are: 1) Inclusion of the target item at the end of the path; 2) maintenance of coherence, ensuring each item in the path is relevant to its adjacent items, thereby enhancing the guiding effectiveness; and 3) utilization of items that users are likely to enjoy, improving user acceptability and keeping users engaged in the interaction session.
Fortunately, Large Language Models (LLMs) possess powerful instruction-following and planning abilities (Shi et al., 2024), demonstrating great potential for generating a coherent influence path that includes target items. In light of this, we propose an LLM-based Influence Path Planning (LLM-IPP) method, which elaborately prompts LLMs to capture user interest shifts and generate coherent and effective influence paths. We design automatic evaluation metrics to compare the performance of LLM-IPP with multiple traditional influence path planning. Moreover, we present various user simulators based on LLMs and traditional recommender models to evaluate the coherence of the influence path and user acceptability towards the influence path.
The contributions of this paper are:
  • To our knowledge, this is the first attempt to integrate LLMs into proactive recommender systems and we design a simple yet effective method for LLMs to achieve influence path planning.
  • To evaluate the performance of LLM-IPP, we design various simulators and metrics to assess user acceptability and the coherence of the influence path.
  • Extensive experiments validate that LLM-IPP significantly surpasses traditional baselines with stronger user acceptability and path coherence.
2. Method
In this study, we propose an innovative method that leverages the instruction-following and path-planning capabilities of LLMs to generate influence paths. Our design involves the inclusion of target items, maintaining coherence, and incorporating items of interest to the user into the instructions, prompting the LLMs to utilize their world knowledge to create paths that meet the conditions. To further enhance the performance, we employ prompt engineering techniques to improve the interest-guiding ability of LLM-IPP.
2.1. Influence Path Generation
LLM-IPP leverages the user’s demographic features and historical data sequence to prompt LLMs to generate the influence path. For example, in the MovieLens-1M dataset, the prompt is shown below:
Prompt for MovieLens-1M dataset
(Optional: ToT) Imagine three different experts answering this question. All experts will write down 1 step of their thinking and then share it with the group. Then all experts will go on to the next step, etc. And show me the result in the end. The question is:
You are a recommender system. Given the user profile and historical data, analyze the user’s interests. Your task is to add no less than five movies between the last movie of historical data and the target movie to connect them. Any adjacent movies should have a strong relation (share the same genre, theme, title, etc.) with each other and make sure the movies are not included in the historical data. The release date of the movies should be before 2001.
(Optional: CoT) Think step by step and make sure.
¡User’s demographic feature¿: Gender/Age/Occupation
¡Historical data¿: sequence of ¡Movie title, Movie genre¿
¡Target item¿: ¡Movie title, Movie genre¿
In the prompt, we set several constraints for the LLMs. (1) The LLMs are required to connect the user interest with the target item, implying that the influence path must include the target item. (2) Direct recommendation of the target item to the user is prohibited; therefore, a minimum length for the influence path is essential. (3) Adjacent items should exhibit a strong relation to instructing the LLMs to generate an influence path with high coherence. Additionally, to make the recommendation more reasonable, the generated items should meet the dataset-specific limitations. For example, for the MovieLens-1M dataset, the items are all movies before 2001, thus we need to constrain the output space in the prompt. For the Last.FM dataset, the items are musical artists before 2011.
We also discovered that the commonly used prompt engineering techniques, such as ToT (Long, 2023) and CoT (Wei et al., 2023), are effective on LLM-IPP. Only minor modifications to these techniques are required to enhance its performance. For example, CoT needs the addition of “Think step by step and make sure.” at the end of the prompt. For ToT, we based on Hubert’s methodology (Hulbert, 2023), employing a straightforward prompting engineering approach to integrate the principal concept of ToT into LLM-IPP.
3. Experiment
In this section, we conduct two experiments. First, we compare the LLM-IPP with other traditional baselines on path planning ability on real-world datasets in the movies and music domain. Second, we compare the LLM-based methods on a cross-domain dataset to compare the performance when using different prompts.
Certain-domain Dataset. We conduct the experiment on two real-world datasets, MovieLens-1M  [^1] https://grouplens.org/datasets/movielens/1m/. and Last.FM  [^2] https://grouplens.org/datasets/hetrec-2011/. dataset. MovieLens-1M dataset contains 6,038 users and 3,883 movies, totaling 5,084 user history item sequences, excluding sequences shorter than 20. Last.FM dataset contains 1,892 users and 17,632 musical artists, totaling 1,297 user history item sequences. In this work, we separate the dataset into training and testing sets accounting for 20% and 80% respectively. The traditional baselines are trained on the training sets and evaluated on the testing sets. LLM-IPP is directly evaluated on the testing sets without training since it is a zero-shot method.
Table 1. Results in MovieLens-1M and Last.FM datasets. We implement five LLM-based user simulators: GPT-3.5, GPT-4, Gemini, Llama2 7b, and Llama3 8b, and one traditional recommender system SASRec to evaluate the acceptability and coherence score. The bold number indicates the best performance, and the underlined number indicates the second-best performance.
Acceptability Coherence SASRec
Method GPT-3.5 GPT-4 Gemini Llama2 Llama3 SASRec GPT-3.5 GPT-4 Gemini Llama2 Llama3 OpenAI $IoI$ $IoR$ $SR$
MovieLens 1M POP 0.409 0.576 0.53 0.601 0.591 400.5 0.483 0.553 0.474 0.530 0.453 0.842 0.801 193.9 0.008
IRS 0.301 0.460 0.42 0.564 0.597 973.3 0.376 0.457 0.489 0.556 0.475 0.800 2.680 651.9 0.268
IRS-fix 0.284 0.438 0.346 0.575 0.580 1010.9 0.365 0.443 0.389 0.540 0.458 0.799 2.680 651.9 1
GPT-3.5 0.629 0.742 0.690 0.622 0.710 736.7 0.682 0.649 0.653 0.590 0.631 0.842 5.409 1192.9 1
GPT-CoT 0.610 0.698 0.679 0.623 0.705 788.9 0.662 0.686 0.652 0.579 0.625 0.841 5.450 1215.6 1
GPT-ToT 0.605 0.682 0.666 0.635 0.718 796.3 0.673 0.701 0.630 0.579 0.617 0.839 5.596 1240.5 1
Last.FM POP 0.334 0.497 0.414 0.526 0.521 2147.5 0.609 0.552 0.518 0.546 0.456 0.810 0.914 619.8 0
IRS 0.406 0.618 0.556 0.567 0.525 1255.4 0.520 0.563 0.492 0.506 0.437 0.811 0.723 516.1 0.142
IRS-fix 0.441 0.563 0.483 0.560 0.524 1627.0 0.541 0.533 0.536 0.525 0.429 0.809 0.723 516.1 1
GPT-3.5 0.532 0.539 0.542 0.582 0.660 4145.6 0.729 0.803 0.652 0.591 0.667 0.802 6.159 1447.5 1
GPT-CoT 0.545 0.558 0.584 0.590 0.672 4002.8 0.725 0.776 0.700 0.589 0.670 0.809 6.354 1544.5 1
GPT-ToT 0.547 0.634 0.618 0.592 0.652 3824.0 0.718 0.731 0.594 0.585 0.657 0.803 6.291 1546.5 1
Cross-domain Dataset. In order to explore the ability of LLM-IPP to guide various item types across diverse data domains, we construct the Cross-domain Dataset. Users’ demographic features are collected from MovieLens-1M and Social circles: Google+ Dataset (Leskovec and Mcauley, 2012). Because the Cross-domain Dataset we constructed lacks the historical item sequence for each user, relying solely on demographic features is insufficient to represent users’ personalities. Furthermore, due to data privacy concerns, most of the user’s demographic features are either vague or not provided. Consequently, we incorporated GPT-3.5 to enrich the demographic features with interests, disinterests, and self-introductions, as demonstrated below. Besides, the target items are sampled from YouTube-8M Dataset (Abu-El-Haija et al., 2016), Wikipedia Article Titles Dataset  [^3] https://www.kaggle.com/datasets/residentmario/wikipedia-article-titles/data., and Twitter Hashtags Dataset (Yang and Leskovec, 2011). Finally, we sample 100 users and 100 target items for LLM-IPP to generate the influence path and demonstrate the performance of LLM-IPP with different prompts within the Cross-domain Dataset.
Prompt for user demographic feature enrichment
Given the demographic features, based on your imagination, expand the user’s demographic features, add a self-introduction, and specify the interests and disinterests in detail.
¡User’s demographic feature¿: Gender/Age/Occupation
3.1. Metrics
To thoroughly evaluate the methods, we employ both LLM-based and traditional metrics. LLM-based metrics utilize independent LLMs to represent the users to provide feedback and evaluate the performance of influence paths. Traditional metrics are based on the simulator implemented by the Transformer model and word embeddings to calculate the performance of influence paths.
3.1.1. LLM-based Simulator
To appropriately evaluate the LLM-based method, we introduce the LLM-based Simulator. Traditional metrics are incapable of revealing the latent relationships between the items, unlike LLMs, due to the lack of world knowledge about the recommended items. The most convincing evaluation method is implementing LLM-IPP on an online social platform and collecting real user feedback, which is significantly expensive. Fortunately, many recent studies on LLMs imply that the LLMs can be an appropriate assessment tool, even for recommender systems (Tokutake and Okamoto, 2024; Zhang et al., 2024). In this paper, we implement commonly used LLMs such as GPT, Gemini, and Llama to provide scores for the influence paths on each metric.
Acceptability (LLMs) measures how much the user accepts the influence path and is interested in the target item by asking the LLMs.
Prompt for acceptability
Given the user profile and the historical data, analyze the user’s interest. Based on this information, would the user be interested in the movies in the influence path step by step? Answer with a probability for each movie between 0 and 1, where 0 means “definitely not interested” and 1 means “definitely interested”. Please explain the reason for each score. If uncertain, make your best guess.
¡User’s demographic feature¿: Gender/Age/Occupation
¡historical data¿: sequence of ¡Movie title, Movie genre¿
¡influence path¿: sequence of ¡Movie title, Movie genre¿
Coherence (LLMs) measures the mean relevance of each of two adjacent items by asking the LLMs.
Prompt for coherence
Given the influence path in the format of A,B,C…, what’s the relevance of each adjacent item? Answer with a score between -1 and 1, where -1 means “definitely not related” and 1 means “definitely related”. Please explain the reason for each score. If uncertain, make your best guess
¡influence path¿: sequence of ¡Movie title, Movie genre¿
3.1.2. Traditional Recommender System Simulator
Except for the LLM-based simulator, we train an independent next-item recommender system that works as a user simulator for evaluation (Bi et al., 2024; Zhu et al., 2023). In this study, SASRec (Kang and McAuley, 2018) is employed as the base model, which utilizes a self-attention-based model that captures both long-term and short-term semantics. SASRec learns user behaviors from the historical item sequence and outputs the probability distribution of items that represent the user’s current preference. In addition, by incorporating word embedding methods and probability theory, we employ the following metrics to evaluate the performance of the influence path.
Success rate ($SR$) measures the ratio of the generated influence path that contains the target item.
Increase of Interest ($IoI$(Zhu et al., 2023) measures the change in the user’s interest in the target item after being persuaded through the influence path.
(1) $IoI=\frac{1}{|U|}\sum_{u=1}^{|U|}(\log{P(i_{t}^{u}|seq^{u}\oplus seq_{t}^{u})}-\log{P(i_{t}^{u}|seq^{u})})$
where $\oplus$ denotes the concatenation of two sequences; $seq^{u}$ is the history item sequence of user $u$; $i_{t}^{u}$ is the target item set for user $u$; $seq_{t}^{u}$ is the generated influence path; $P(i|seq)$ is the probability of item $i$ when input the sequence $seq$ into the SASRec model.
Increase of Rank ($IoR$(Zhu et al., 2023) measures the increase of the ranking for the target item after being persuaded through the influence path.
(2) $IoR=-\frac{1}{|U|}\sum_{u=1}^{|U|}(R(i_{t}^{u}|seq^{u}\oplus seq_{t}^{u})-R(i_{t}^{u}|seq^{u}))$
where $R(i|seq)$ is the rank of item $i$ when input the sequence $seq$ into the SASRec model.
Acceptability (SASRec) measures user acceptability by calculating the mean ranking of the items in the influence path.
(3) $Acceptability=\frac{1}{|U|}\sum_{u=1}^{|U|}\frac{1}{|seq^{u}_{t}|}\sum_{i\in seq_{t}^{u}}R(i|seq^{u})$
Coherence (OpenAI) measures the relevance of each of two adjacent items in the influence path by calculating the mean cosine similarity of OpenAI Embeddings (Neelakantan et al., 2022).
(4) $Coherence=\frac{1}{|U|}\sum_{u=1}^{|U|}\frac{1}{|seq^{u}_{t}|}\sum_{i_{j},i_{j+1}\in seq^{u}_{t}}cos\_sim(i_{j},i_{j+1})$
3.2. Baselines & Proposed Methods
We implement three baselines and three LLM-IPP methods.
IRS (Zhu et al., 2023) is the SOTA traditional proactive recommender system. Improve the performance by designing a Transformer-based sequential model.
IRS-fix fixes the baseline’s low success rate. We append the target item to the IRS-generated influence path if it doesn’t contain the target item to reach a 100% success rate.
POP sorts all the items by occurrence and recommends the most popular items to the user.
LLM-IPP (GPT-3.5) is the base version of LLM-IPP.
LLM-IPP (GPT-CoT) enables complex reasoning capabilities via intermediate reasoning steps.
LLM-IPP (GPT-ToT) introduces a framework that generalizes over CoT prompting and encourages exploration over thoughts that serve as intermediate steps for general problem-solving with language models.
Table 2. Results in Cross-domain Dataset.
Acceptability Coherence
Method GPT-3.5 GPT-4 Gemini OpenAI GPT-4
GPT-3.5 0.65 0.63 0.64 0.08 0.65
GPT-CoT 0.73 0.75 0.58 0.11 0.70
GPT-ToT 0.84 0.77 0.55 0.19 0.81
3.3. Analysis
Table 1 presents the results of the methods evaluated in the MovieLens-1M dataset and the Last.FM dataset. From the table, it is evident that LLM-based methods achieve higher scores on most of the metrics than other baselines. For the traditional metrics, all the LLM-based methods significantly outperform traditional baselines in $IoI$ and $IoR$, demonstrating that LLM-IPP effectively improves the user’s interest towards the target item. For the LLM-based metrics, an interesting trade-off between the influence path’s acceptability and coherence is observed, especially in Last.FM dataset. This phenomenon is possible because when the items in the influence path are closer to the user’s interests and achieve a higher acceptability score, it will result in the items not being evenly distributed along the influence path and consequently cause a lower coherence score.
However, the results from the Cross-domain Dataset, as shown in Table 2, are different. The performance of GPT-ToT surpasses that of other methods in all metrics without the trade-off phenomenon that appeared in the Certain-domain Dataset. This can be attributed to the fact that the LLMs in the Cross-domain Dataset have richer user features and face fewer domain limitations than those in the MovieLens-1M and Last.FM datasets. These facts prove that LLM-based models can achieve a higher upper limit, and prompt engineering techniques considerably improve the path planning ability on the Cross-domain Dataset.
Case study: In addition to the quantitative study, we conduct a real case to demonstrate the effectiveness of LLM-IPP. As shown in Table 3, The Great Dictator is selected as the target item. At first, LLM-IPP recommends movies in the genre of “Drama”. After a few attempts, Being There, with the genres of both “Drama” and “Comedy,” is recommended. Then finally to the target item.
However, LLM-IPP reaches further beyond the movie names and genres, which are the only details provided by the MovieLens-1M dataset. With LLM’s great ability in natural language understanding and world knowledge, LLM-IPP knows the movies’ cast, director, content, and other movie attributes. In this example, based on the user’s other viewing history, LLM-IPP considers this user is likely to enjoy movies with the genres of “War” and “Drama” and enjoys movies with strong narratives and suspenseful plots. According to the recommendation reasons generated by LLM-IPP, Path of Glory and Dr. Strangelove are both directed by Stanley Kubrick and have a war theme. Network and Being There are both about media critique and dark comedy. These understandings about the movies help to build a more practical influence path and reach higher acceptability and coherence scores. The results affirm the effectiveness of LLM-IPP in smoothly leading the user preference to the target items.
Table 3. An example of an influence path.
Name Genre
The last movie in the viewing history
Frequency (2000) Drama, Thriller
Influence path
Paths of Glory (1957) Drama, War
Dr. Strangelove (1964) Comedy, War
Network (1976) Drama
Being There (1979) Comedy, Drama
To Be or Not to Be (1942) Comedy, Romance, War
The Great Dictator (1940) Comedy
4. Conclusion & Discussion
This paper proposed an LLM-based Influence Path Planning (LLM-IPP) method for proactive recommendation, which can generate the influence path with high coherence and user acceptability. We explored several prompting techniques to identify the optimal prompt to instruct LLMs for influence path planning. Additionally, to evaluate LLM-IPP, we implemented various LLM-based and traditional simulators and metrics to ensure a comprehensive assessment of user acceptability and path coherence. Experimental results demonstrate that the LLM-IPP outperforms traditional proactive recommender systems in terms of path coherence and user acceptability.
As the first work to apply LLMs to proactive recommender systems, we summarize the future work in the following directions: (1) While extensive user studies are essential, it is hard for participants to simulate the interest shifts across multi-round recommendations. How to design an effective user study for proactive recommendation is a future direction. (2) It is feasible to further fine-tune LLM-IPP with real-world data, injecting real-world user interest shifting patterns into LLMs for influence path planning.

### v2

@@
-Incorporate LLMs with Influential Recommender System
+Leveraging LLMs for Influence Path Planning in Proactive Recommendation
 Abstract.
-Recommender systems have achieved increasing accuracy over the years.
-However, this precision often leads users to narrow their interests, resulting in issues such as limited diversity and the creation of echo chambers.
-Current research addresses these challenges through proactive recommender systems by recommending a sequence of items (called influence path) to guide user interest to the target item.
-However, existing methods struggle to construct a coherent influence path that build up with items the user is likely to enjoy.
-In this paper, we leverage the Large Language Model’s (LLMs) exceptional ability for path planning and instruction following, introducing a novel approach named LLM-based Influence Path Planing (LLM-IPP).
-Our approach maintains coherence between consecutive recommendations, and enhances user acceptability to the recommended items.
-To evaluate LLM-IPP, we implement various user simulators and metrics to measure user acceptability and path coherence.
-Experimental results demonstrate that LLM-IPP significantly outperforms traditional proactive recommender systems.
-This study is pioneering in integrating LLMs into proactive recommender systems, offering a reliable and user-engaging methodology for future recommendation technologies.
+Recommender systems are pivotal in Internet social platforms, yet they often cater to users’ historical interests, leading to critical issues like echo chambers.
+To broaden user horizons, proactive recommender systems aim to guide user interest to gradually like a target item beyond historical interests through an influence path,i.e., a sequence of recommended items.
+As a representative, Influential Recommender System (IRS) designs a sequential model for influence path planning but faces issues of lacking target item inclusion and path coherence.
+To address the issues, we leverage the advanced planning capabilities of Large Language Models (LLMs) and propose an LLM-based Influence Path Planning (LLM-IPP) method.
+LLM-IPP generates coherent and effective influence paths by capturing user interest shifts and item characteristics.
+We introduce novel evaluation metrics and user simulators to benchmark LLM-IPP against traditional methods.
+Our experiments demonstrate that LLM-IPP significantly enhances user acceptability and path coherence, outperforming existing approaches.
 Proactive Recommender Systems, Large Language Model, Influence Path, User Simulator
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXbooktitle: ccs: Information systems Recommender systems
+journalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715537isbn: 979-8-4007-1331-6/2025/04ccs: Information systems Recommender systems
 1. Introduction
 Figure 1. Illustration of Passive and Proactive Recommendation.
-Recommender systems play an important role in enhancing user experiences.
-Traditionally, recommender systems infer user interest from historical behavior patterns and recommend items aligned with the user’s interest.
-However, it will narrow the user’s interest and cause the echo chamber phenomenon by continually catering to historical interest (Gao et al., 2023b, a).
-To avoid this problem, researchers have proposed the task of proactive recommender system (Zhu et al., 2023), which aims to actively guide the user interest with a sequence of items (called influence path) toward a given target item, as shown in Figure 1.
-The influence path gradually guides shifts in user interests through continuous recommendations, facilitating the discovery of new interests.
-Therefore, influence path planning is crucial for enhancing the efficacy of proactive recommender systems in guiding user interests.
-The most pertinent work is Influential Recommender System (IRS) (Zhu et al., 2023), which frames proactive recommendation as a sequential recommendation problem and constructs the influence path using a Transformer-based sequential model.
+Recommender system plays an important role in Internet social platforms.
+Traditionally, recommender systems infer user interest from historical behavior patterns and recommend items aligned with the user’s interest (Sun et al., 2020; Pandey et al., 2024; Mao et al., 2024; Shi et al., 2024).
+However, it will narrow the user’s interest and cause the echo chamber phenomenon by continually catering to historical interest (Gao et al., 2023b, a, 2025; Chen et al., 2025).
+To avoid this problem, researchers have proposed the task of proactive recommender system (Zhu et al., 2023), which aims to actively guide the user interest with a sequence of items (called influence path) toward a given target item beyond historical interests, as shown in Figure 1.
+The influence path gradually guides shifts in user interests through continuous recommendation exposure, facilitating the discovery of new interests.
+As such, influence path planning is the most crucial step for enhancing the efficacy of proactive recommender systems.
+The most pertinent work is Influential Recommender System (IRS) (Zhu et al., 2023), which frames proactive recommendation as a sequential recommendation problem and constructs the influence path using a Transformer-based sequential model (Vaswani et al., 2017).
 Such a sequential model learns the interest-shifting patterns from users’ historical data and encodes the target item’s characteristics to generate the influence path.
-However, IRS encounters several critical issues: 1) There is a low probability of the target item being included in the influence path, with a success rate below 30% on MovieLens-1M (refer to Table 1); and 2) the adjacent items in the path are not sufficiently coherent, hindering the effective guidance of user interest.
+However, IRS encounters several critical issues: 1) There is a low probability of target item inclusion in the influence path, with a success rate below 30% on MovieLens-1M (refer to Table 1); and 2) the adjacent items in the path are not sufficiently coherent, hindering the effective guidance of user interest.
 Moreover, directly incorporating the target item into the influence path to address the first issue will further diminish the path’s coherence.
-Inspired by these, the key objectives of influence path planning are: 1) Inclusion of the target item at the end of the path; 2) maintenance of coherence, ensuring each item in the path is relevant to its adjacent items, thereby enhancing the guiding effectiveness; and 3) utilization of items that users are likely to enjoy, improving user acceptability and keeping users engaged in the interaction session.
-Fortunately, Large Language Models (LLMs) possess powerful instruction-following and planning abilities (Shi et al., 2024), demonstrating great potential for generating a coherent influence path that includes target items.
-In light of this, we propose an LLM-based Influence Path Planning (LLM-IPP) method, which elaborately prompts LLMs to capture user interest shifts and generate coherent and effective influence paths.
-We design automatic evaluation metrics to compare the performance of LLM-IPP with multiple traditional influence path planning.
+In light of these, the key objectives of influence path planning are: 1) incorporating the target item at the end of the path; 2) improving path coherence for guiding smoothness and effectiveness by ensuring adjacent items in the path are relevant; and 3) using items that users are likely to enjoy for guidance, improving user acceptability and keeping users engaged in the interaction session.
+Fortunately, Large Language Models (LLMs) possess powerful instruction-following and planning abilities (Shi et al., 2024; Yang et al., 2024; Ren et al., 2024), demonstrating great potential for generating a coherent influence path that includes target items.
+In this light, we propose an LLM-based Influence Path Planning (LLM-IPP) method, which elaborately prompts LLMs to capture user interest shifts and generate coherent and effective influence paths through prompt engineering techniques.
+We further design automatic evaluation metrics to compare the performance of LLM-IPP with multiple traditional influence path planning methods.
 Moreover, we present various user simulators based on LLMs and traditional recommender models to evaluate the coherence of the influence path and user acceptability towards the influence path.
+Empirical results validate the superiority of LLM-IPP in terms of path coherence and user acceptability.
+We release our code and data at https://github.com/gnaWeinrE/LLM-IPP.
 The contributions of this paper are:
-  • To our knowledge, this is the first attempt to integrate LLMs into proactive recommender systems and we design a simple yet effective method for LLMs to achieve influence path planning.
-  • To evaluate the performance of LLM-IPP, we design various simulators and metrics to assess user acceptability and the coherence of the influence path.
+  • To our knowledge, this is the first attempt to integrate LLMs into proactive recommendation and we design a simple yet effective method for LLMs to achieve influence path planning.
+  • To evaluate the performance of LLM-IPP, we design diverse user simulators and metrics to assess user acceptability and the coherence of the influence path.
   • Extensive experiments validate that LLM-IPP significantly surpasses traditional baselines with stronger user acceptability and path coherence.
-2. Method
-In this study, we propose an innovative method that leverages the instruction-following and path-planning capabilities of LLMs to generate influence paths.
-Our design involves the inclusion of target items, maintaining coherence, and incorporating items of interest to the user into the instructions, prompting the LLMs to utilize their world knowledge to create paths that meet the conditions.
-To further enhance the performance, we employ prompt engineering techniques to improve the interest-guiding ability of LLM-IPP.
-2.1. Influence Path Generation
-LLM-IPP leverages the user’s demographic features and historical data sequence to prompt LLMs to generate the influence path. For example, in the MovieLens-1M dataset, the prompt is shown below:
-Prompt for MovieLens-1M dataset
-(Optional: ToT) Imagine three different experts answering this question.
-All experts will write down 1 step of their thinking and then share it with the group.
-Then all experts will go on to the next step, etc.
-And show me the result in the end.
-The question is:
-You are a recommender system.
-Given the user profile and historical data, analyze the user’s interests.
-Your task is to add no less than five movies between the last movie of historical data and the target movie to connect them.
-Any adjacent movies should have a strong relation (share the same genre, theme, title, etc.) with each other and make sure the movies are not included in the historical data.
-The release date of the movies should be before 2001.
-(Optional: CoT) Think step by step and make sure.
-¡User’s demographic feature¿: Gender/Age/Occupation
-¡Historical data¿: sequence of ¡Movie title, Movie genre¿
-¡Target item¿: ¡Movie title, Movie genre¿
-In the prompt, we set several constraints for the LLMs.
-(1) The LLMs are required to connect the user interest with the target item, implying that the influence path must include the target item.
-(2) Direct recommendation of the target item to the user is prohibited; therefore, a minimum length for the influence path is essential.
-(3) Adjacent items should exhibit a strong relation to instructing the LLMs to generate an influence path with high coherence.
-Additionally, to make the recommendation more reasonable, the generated items should meet the dataset-specific limitations.
-For example, for the MovieLens-1M dataset, the items are all movies before 2001, thus we need to constrain the output space in the prompt.
-For the Last.FM dataset, the items are musical artists before 2011.
-We also discovered that the commonly used prompt engineering techniques, such as ToT (Long, 2023) and CoT (Wei et al., 2023), are effective on LLM-IPP.
-Only minor modifications to these techniques are required to enhance its performance.
-For example, CoT needs the addition of “Think step by step and make sure.” at the end of the prompt.
-For ToT, we based on Hubert’s methodology (Hulbert, 2023), employing a straightforward prompting engineering approach to integrate the principal concept of ToT into LLM-IPP.
-3. Experiment
-In this section, we conduct two experiments.
-First, we compare the LLM-IPP with other traditional baselines on path planning ability on real-world datasets in the movies and music domain.
-Second, we compare the LLM-based methods on a cross-domain dataset to compare the performance when using different prompts.
-Certain-domain Dataset.
-We conduct the experiment on two real-world datasets, MovieLens-1M  [^1] https://grouplens.org/datasets/movielens/1m/.
-and Last.FM  [^2] https://grouplens.org/datasets/hetrec-2011/.
-dataset.
-MovieLens-1M dataset contains 6,038 users and 3,883 movies, totaling 5,084 user history item sequences, excluding sequences shorter than 20.
-Last.FM dataset contains 1,892 users and 17,632 musical artists, totaling 1,297 user history item sequences.
-In this work, we separate the dataset into training and testing sets accounting for 20% and 80% respectively.
-The traditional baselines are trained on the training sets and evaluated on the testing sets.
-LLM-IPP is directly evaluated on the testing sets without training since it is a zero-shot method.
 Table 1.
@@
 GPT-ToT 0.547 0.634 0.618 0.592 0.652 3824.0 0.718 0.731 0.594 0.585 0.657 0.803 6.291 1546.5 1
-Cross-domain Dataset.
-In order to explore the ability of LLM-IPP to guide various item types across diverse data domains, we construct the Cross-domain Dataset.
-Users’ demographic features are collected from MovieLens-1M and Social circles: Google+ Dataset (Leskovec and Mcauley, 2012).
-Because the Cross-domain Dataset we constructed lacks the historical item sequence for each user, relying solely on demographic features is insufficient to represent users’ personalities.
-Furthermore, due to data privacy concerns, most of the user’s demographic features are either vague or not provided.
-Consequently, we incorporated GPT-3.5 to enrich the demographic features with interests, disinterests, and self-introductions, as demonstrated below.
-Besides, the target items are sampled from YouTube-8M Dataset (Abu-El-Haija et al., 2016), Wikipedia Article Titles Dataset  [^3] https://www.kaggle.com/datasets/residentmario/wikipedia-article-titles/data., and Twitter Hashtags Dataset (Yang and Leskovec, 2011).
-Finally, we sample 100 users and 100 target items for LLM-IPP to generate the influence path and demonstrate the performance of LLM-IPP with different prompts within the Cross-domain Dataset.
-Prompt for user demographic feature enrichment
-Given the demographic features, based on your imagination, expand the user’s demographic features, add a self-introduction, and specify the interests and disinterests in detail.
-¡User’s demographic feature¿: Gender/Age/Occupation
+2. Method
+In this study, we propose an innovative method that leverages the instruction-following and path-planning capabilities of LLMs to generate influence paths.
+Our design involves the inclusion of target items, maintaining coherence, and incorporating items of interest to the user into the instructions, prompting the LLMs to utilize their world knowledge to create paths that meet the conditions.
+To further enhance the performance, we employ prompt engineering techniques to improve the interest-guiding ability of LLM-IPP.
+2.1. Influence Path Generation
+LLM-IPP leverages the user’s demographic features and historical data sequence to prompt LLMs to generate the influence path.
+For example, in the MovieLens-1M dataset, the prompt is as shown on the right.
+In the prompt, we set several constraints for the LLMs.
+(1) The LLMs are required to connect the user interest with the target item, implying that the influence path must include the target item.
+(2) Direct recommendation of the target item to the user is prohibited; therefore, a minimum length for the influence path is essential.
+(3) Adjacent items should exhibit a strong relation to instructing the LLMs to generate an influence path with high coherence.
+Additionally, to make the recommendation more reasonable, the generated items should meet the dataset-specific limitations.
+For example, for the MovieLens-1M dataset, the items are all movies before 2001, thus we need to constrain the output space in the prompt.
+For the Last.FM dataset, the items are musical artists before 2011.
+Prompt for MovieLens-1M dataset (Optional: ToT) Imagine three different experts answering this question.
+All experts will write down 1 step of their thinking and then share it with the group.
+Then all experts will go on to the next step, etc.
+And show me the result in the end.
+The question is: You are a recommender system.
+Given the user profile and historical data, analyze the user’s interests.
+Your task is to add no less than five movies between the last movie of historical data and the target movie to connect them.
+Any adjacent movies should have a strong relation (share the same genre, theme, title, etc.) with each other and make sure the movies are not included in the historical data.
+The release date of the movies should be before 2001.
+(Optional: CoT) Think step by step and make sure.
+¡User’s demographic feature¿: Gender/Age/Occupation ¡Historical data¿: sequence of ¡Movie title, Movie genre¿ ¡Target item¿: ¡Movie title, Movie genre¿
+We also discovered that the commonly used prompt engineering techniques, such as ToT (Long, 2023) and CoT (Wei et al., 2022), are effective on LLM-IPP.
+Only minor modifications to these techniques are required to enhance its performance.
+For example, CoT needs the addition of “Think step by step and make sure.” at the end of the prompt.
+For ToT, we based on Hubert’s methodology (Hulbert, 2023), employing a straightforward prompting engineering approach to integrate the principal concept of ToT into LLM-IPP.
+3. Experiment
+In this section, we conduct experiments on two real-world datasets and design simulators with corresponding metrics to evaluate the influence paths.
+LLMs are used to simulate users to provide feedback and leverage LLM-based metrics to evaluate the performance of influence paths.
+Traditional recommender models can also be developed as user simulators, usually implemented by the Transformer model and word embeddings (Kang and McAuley, 2018) to assess the influence paths.
+Certain-domain Dataset.
+We conduct the experiment on two real-world datasets, MovieLens-1M  [^1] https://grouplens.org/datasets/movielens/1m/.
+and Last.FM  [^2] https://grouplens.org/datasets/hetrec-2011/.
+dataset.
+MovieLens-1M dataset contains 6,038 users and 3,883 movies, totaling 5,084 user history item sequences, excluding sequences shorter than 20.
+Last.FM dataset contains 1,892 users and 17,632 musical artists, totaling 1,297 user history item sequences.
+In this work, the target items are randomly sampled from all items.
+For training, we separate the dataset into training and testing sets accounting for 20% and 80% respectively.
+The traditional baselines are trained on the training sets and evaluated on the testing sets.
+LLM-IPP is directly evaluated on the testing sets without training since it is a zero-shot method.
 3.1. Metrics
@@
 Traditional metrics are based on the simulator implemented by the Transformer model and word embeddings to calculate the performance of influence paths.
+3.1.1. Traditional Recommender System Simulator
+We train an independent next-item recommender system that works as a user simulator for evaluation (Bi et al., 2024; Zhu et al., 2023).
+In this study, SASRec (Kang and McAuley, 2018) is employed as the base model, which utilizes a self-attention-based model that captures both long-term and short-term semantics.
+SASRec learns user behaviors from the historical item sequence and outputs the probability distribution of items that represent the user’s current preference.
+In addition, by incorporating word embedding methods and probability theory, we employ the following metrics to evaluate the performance of the influence path.
+Success rate (SR) measures the ratio of the generated influence path that contains the target item.
+Increase of Interest (IoI) (Zhu et al., 2023) measures the change in the user’s interest in the target item after being persuaded through the influence path.
+(1) $\textrm{IoI}=\frac{1}{|U|}\sum_{u=1}^{|U|}(\log{P(i_{t}^{u}|seq^{u}\oplus seq_{t}^{u})}-\log{P(i_{t}^{u}|seq^{u})}),$
+where $\oplus$ denotes the concatenation of two sequences; $seq^{u}$ is the history item sequence of user $u$; $i_{t}^{u}$ is the target item set for user $u$; $seq_{t}^{u}$ is the generated influence path; $P(i|seq)$ is the probability of item $i$ when input the sequence $seq$ into the SASRec model.
+Increase of Rank (IoR) (Zhu et al., 2023) measures how much the target item’s ranking improves after persuasion via the influence path.
+(2) $\textrm{IoR}=-\frac{1}{|U|}\sum_{u=1}^{|U|}(R(i_{t}^{u}|seq^{u}\oplus seq_{t}^{u})-R(i_{t}^{u}|seq^{u})),$
+where $R(i|seq)$ is the rank of item $i$ when input the sequence $seq$ into the SASRec model.
+Acceptability (SASRec) measures user acceptability by calculating the mean ranking of the items in the influence path.
+(3) $\textrm{Acceptability}=\frac{1}{|U|}\sum_{u=1}^{|U|}\frac{1}{|seq^{u}_{t}|}\sum_{i\in seq_{t}^{u}}R(i|seq^{u}).$
+Coherence (OpenAI) measures the relevance of each of two adjacent items in the influence path by calculating the mean cosine similarity of OpenAI Embeddings (Neelakantan et al., 2022).
+(4) $\textrm{Coherence}=\frac{1}{|U|}\sum_{u=1}^{|U|}\frac{1}{|seq^{u}_{t}|}\sum_{i_{j},i_{j+1}\in seq^{u}_{t}}\textrm{cos\_sim}(i_{j},i_{j+1}).$
 3.1.2. LLM-based Simulator
-To appropriately evaluate the LLM-based method, we introduce the LLM-based Simulator.
+Except for the traditional recommender system simulator, to appropriately evaluate the LLM-based method, we introduce the LLM-based Simulator.
 Traditional metrics are incapable of revealing the latent relationships between the items, unlike LLMs, due to the lack of world knowledge about the recommended items.
@@
 In this paper, we implement commonly used LLMs such as GPT, Gemini, and Llama to provide scores for the influence paths on each metric.
-Acceptability (LLMs) measures how much the user accepts the influence path and is interested in the target item by asking the LLMs.
-Prompt for acceptability
-Given the user profile and the historical data, analyze the user’s interest.
+Acceptability (LLMs) measures the user’s acceptance of the influence path and interest in the target item via LLMs.
+Prompt for measuring user acceptability Given the user profile and the historical data, analyze the user’s interest.
 Based on this information, would the user be interested in the movies in the influence path step by step?
@@
 If uncertain, make your best guess.
-¡User’s demographic feature¿: Gender/Age/Occupation
-¡historical data¿: sequence of ¡Movie title, Movie genre¿
-¡influence path¿: sequence of ¡Movie title, Movie genre¿
+¡User’s demographic feature¿: Gender/Age/Occupation ¡historical data¿: sequence of ¡Movie title, Movie genre¿ ¡influence path¿: sequence of ¡Movie title, Movie genre¿
 Coherence (LLMs) measures the mean relevance of each of two adjacent items by asking the LLMs.
-Prompt for coherence
-Given the influence path in the format of A,B,C…, what’s the relevance of each adjacent item?
+Prompt for measuring path coherence Given the influence path in the format of A,B,C…, what’s the relevance of each adjacent item?
 Answer with a score between -1 and 1, where -1 means “definitely not related” and 1 means “definitely related”.
 Please explain the reason for each score.
-If uncertain, make your best guess
-¡influence path¿: sequence of ¡Movie title, Movie genre¿
-3.1.2. Traditional Recommender System Simulator
-Except for the LLM-based simulator, we train an independent next-item recommender system that works as a user simulator for evaluation (Bi et al., 2024; Zhu et al., 2023).
-In this study, SASRec (Kang and McAuley, 2018) is employed as the base model, which utilizes a self-attention-based model that captures both long-term and short-term semantics.
-SASRec learns user behaviors from the historical item sequence and outputs the probability distribution of items that represent the user’s current preference.
-In addition, by incorporating word embedding methods and probability theory, we employ the following metrics to evaluate the performance of the influence path.
-Success rate ($SR$) measures the ratio of the generated influence path that contains the target item.
-Increase of Interest ($IoI$(Zhu et al., 2023) measures the change in the user’s interest in the target item after being persuaded through the influence path.
-(1) $IoI=\frac{1}{|U|}\sum_{u=1}^{|U|}(\log{P(i_{t}^{u}|seq^{u}\oplus seq_{t}^{u})}-\log{P(i_{t}^{u}|seq^{u})})$
-where $\oplus$ denotes the concatenation of two sequences; $seq^{u}$ is the history item sequence of user $u$; $i_{t}^{u}$ is the target item set for user $u$; $seq_{t}^{u}$ is the generated influence path; $P(i|seq)$ is the probability of item $i$ when input the sequence $seq$ into the SASRec model.
-Increase of Rank ($IoR$(Zhu et al., 2023) measures the increase of the ranking for the target item after being persuaded through the influence path.
-(2) $IoR=-\frac{1}{|U|}\sum_{u=1}^{|U|}(R(i_{t}^{u}|seq^{u}\oplus seq_{t}^{u})-R(i_{t}^{u}|seq^{u}))$
-where $R(i|seq)$ is the rank of item $i$ when input the sequence $seq$ into the SASRec model.
-Acceptability (SASRec) measures user acceptability by calculating the mean ranking of the items in the influence path.
-(3) $Acceptability=\frac{1}{|U|}\sum_{u=1}^{|U|}\frac{1}{|seq^{u}_{t}|}\sum_{i\in seq_{t}^{u}}R(i|seq^{u})$
-Coherence (OpenAI) measures the relevance of each of two adjacent items in the influence path by calculating the mean cosine similarity of OpenAI Embeddings (Neelakantan et al., 2022).
-(4) $Coherence=\frac{1}{|U|}\sum_{u=1}^{|U|}\frac{1}{|seq^{u}_{t}|}\sum_{i_{j},i_{j+1}\in seq^{u}_{t}}cos\_sim(i_{j},i_{j+1})$
+If uncertain, make your best guess ¡influence path¿: sequence of ¡Movie title, Movie genre¿
 3.2. Baselines & Proposed Methods
@@
 LLM-IPP (GPT-ToT) introduces a framework that generalizes over CoT prompting and encourages exploration over thoughts that serve as intermediate steps for general problem-solving with language models.
-Table 2. Results in Cross-domain Dataset.
-Acceptability Coherence
-Method GPT-3.5 GPT-4 Gemini OpenAI GPT-4
-GPT-3.5 0.65 0.63 0.64 0.08 0.65
-GPT-CoT 0.73 0.75 0.58 0.11 0.70
-GPT-ToT 0.84 0.77 0.55 0.19 0.81
 3.3. Analysis
@@
 This phenomenon is possible because when the items in the influence path are closer to the user’s interests and achieve a higher acceptability score, it will result in the items not being evenly distributed along the influence path and consequently cause a lower coherence score.
-However, the results from the Cross-domain Dataset, as shown in Table 2, are different.
-The performance of GPT-ToT surpasses that of other methods in all metrics without the trade-off phenomenon that appeared in the Certain-domain Dataset.
-This can be attributed to the fact that the LLMs in the Cross-domain Dataset have richer user features and face fewer domain limitations than those in the MovieLens-1M and Last.FM datasets.
-These facts prove that LLM-based models can achieve a higher upper limit, and prompt engineering techniques considerably improve the path planning ability on the Cross-domain Dataset.
-Case study: In addition to the quantitative study, we conduct a real case to demonstrate the effectiveness of LLM-IPP.
+Case study.
+In addition to the quantitative study, we conduct a real case to demonstrate the effectiveness of LLM-IPP.
 As shown in Table 2, The Great Dictator is selected as the target item.
@@
 The Great Dictator (1940) Comedy
+User Study.
+We conduct a double-blind human test to evaluate the performance of LLM-IPP.
+We recruit three volunteers who are experienced in the field of information retrieval.
+For the evaluation, we create a total of 100 examples.
+For each example, we provide the volunteers with the user’s historical data along with two influence paths generated by the baseline model (IRS) and LLM-IPP (GPT-ToT), respectively.
+The volunteers are asked to determine the superior influence path.
+The preference rates for LLM-IPP among the three volunteers were 66.7%, 63.6%, and 84.2%.
+The average Cohen’s kappa score was 0.63, indicating a high level of agreement among the volunteers.
 4. Conclusion & Discussion
@@
 Experimental results demonstrate that the LLM-IPP outperforms traditional proactive recommender systems in terms of path coherence and user acceptability.
-As the first work to apply LLMs to proactive recommender systems, we summarize the future work in the following directions: (1) While extensive user studies are essential, it is hard for participants to simulate the interest shifts across multi-round recommendations.
-How to design an effective user study for proactive recommendation is a future direction.
-(2) It is feasible to further fine-tune LLM-IPP with real-world data, injecting real-world user interest shifting patterns into LLMs for influence path planning.
+As the first work to apply LLMs to proactive recommender systems, we leave many promising research directions: (1) While extensive user studies are essential, it is hard for participants to simulate the interest shifts across multi-round recommendations.
+How to design an effective user evaluation for proactive recommendation is a future direction.
+(2) It is feasible to fine-tune LLM-IPP with real-world data, injecting real-world user interest shifting patterns into LLMs for influence path planning.
+(3) In addition to prompt engineering, other optimization techniques related to LLMs can be applied to proactive recommendation systems.
+For instance, employing LLM-based agents that use memory mechanisms to improve user satisfaction or fine-tuning LLMs to better align them with proactive recommendation tasks may enhance their performance.
+Acknowledgements
+This work is supported by the National Natural Science Foundation of China (62402470,62272437,U21B2026), the Fundamental Research Funds for the Central Universities of China (WK2100000053, PA2024GDSK0107), Anhui Provincial Natural Science Foundation (2408085QF189), and the Postdoctoral Fellowship Program of CPSF (GZC20241643).
+This research is supported by the advanced computing resources provided by the Supercomputing Center of the USTC.